### v1

# Read Between the Layers: Leveraging Intra-Layer Representations for Rehearsal-Free Continual Learning with Pre-Trained Models

Abstract
We address the Continual Learning (CL) problem, where a model has to learn a sequence of tasks from non-stationary distributions while preserving prior knowledge as it encounters new experiences. With the advancement of foundation models, CL research has shifted focus from the initial learning-from-scratch paradigm to the use of generic features from large-scale pre-training. However, existing approaches to CL with pre-trained models only focus on separating the class-specific features from the final representation layer and neglect the power of intermediate representations that capture low- and mid-level features naturally more invariant to domain shifts. In this work, we propose LayUP, a new class-prototype-based approach to continual learning that leverages second-order feature statistics from multiple intermediate layers of a pre-trained network. Our method is conceptually simple, does not require any replay buffer, and works out of the box with any foundation model. LayUP improves over the state-of-the-art on four of the seven class-incremental learning settings at a considerably reduced memory and computational footprint compared with the next best baseline. Our results demonstrate that fully exhausting the representational capacities of pre-trained models in CL goes far beyond their final embeddings. The code will be made publicly available upon acceptance.
**footnotetext: Equal contribution.
1 Introduction
Figure 1: Overview of LayUP: Enriching last layer representations of pre-trained models with features from intermediate layers leads to better calibrated similarity measures for prototype-based classification and more robustness to domain gaps in the continual learning setting.
Continual Learning (CL) is a subfield of machine learning that focuses on enabling models to learn from a stream of data and adapt to novel concepts without forgetting previously acquired knowledge [45, 6]. While traditional works on CL primarily focus on the learning-from-scratch paradigm, the introduction of large foundation models has initiated a growing interest in developing CL methods upon the powerful representations resulting from large-scale pre-training. Continual learning with pre-trained models builds on the assumption that a large all-purpose feature extractor provides strong knowledge transfer capabilities and great robustness to catastrophic forgetting [39] during incremental adaptation to downstream tasks.
Such methods for downstream adaptation can be categorized into two general strategies [61, 40]: (i) Carefully fine-tuning the pre-trained representations (full body adaptation), or (ii) keeping the pre-trained backbone fixed while learning additional parameters (such as prompts [31, 35, 22], adapters [50, 19, 20, 5, 7], or other methods [25, 38, 34]). Both strategies have in common that they do not rely on rehearsal from past data to perform well, unlike many of the top-performing CL strategies used for models trained from scratch. Despite their merits, most approaches to CL with pre-trained models that fall under one of those strategies update a fully connected classification head (linear probe) via iterative gradient methods during training and inference, which is subject to task-recency bias [37, 54, 62] and catastrophic forgetting [69, 48]. Class-prototype methods [72, 21, 40] have been established as a viable alternative, as they average extracted features directly, which is not only computationally efficient but also superior to CL methods applied to the trainable linear probe.
Despite their effectiveness, contemporary class-prototype methods for CL use only the features obtained by the last layer of the backbone for class-prototype construction. However, as the discrepancy between the pre-training and fine-tuning domain increases, the high-level last-layer features might not generalize well enough to provide a good class separability in the target domain anymore. The challenge is therefore to find a prototype-based classifier that leverages the representations of a pre-trained feature extractor holistically, ideally at comparably low memory and computational cost.
In this work, we hypothesize that representations are most effectively leveraged from multiple layers to construct a classifier based on first-order (class-prototypes) and second-order (Gram matrix) feature statistics. Such a strategy is inspired by works on Neural Style Transfer [13, 23], which involve applying the artistic style of one image to the content of another image, a problem that can be viewed through the lens of domain adaptation [32]. To achieve neural style transfer, a model has to disentangle image information into content and style (e.g., textures, patterns, colors). While content information is expressed in terms of activations of features at different layers (which translates to calculating intra-layer class-prototypes in our case), style information is expressed as correlations between features of different layers (which corresponds to us calculating intra-layer Gram matrices).
We thus propose Intra-Layer Universal Prototypes (LayUP), a class-prototype method for CL that is based on the aforementioned strategy to disentangle style and content information of images at multiple layers of a pre-trained network. An overview of our method is given in Fig. 1. LayUP obtains rich and fine-grained information about images to perform a more informed classification that is less sensitive to domain shifts. We additionally experiment with different parameter-efficient fine-tuning strategies to further refine the intra-layer representations obtained by LayUP. Across several CL benchmarks and learning settings, our method provides improvements and reduces the gap to upper bound performance by up to 67% (Stanford Cars-196 in the CIL setting) compared with the next best baseline while reducing its memory and compute requirements by 81% and up to 90%, respectively.
Our contributions are threefold: (1) We present and examine in detail the CL strategy with pre-trained models and show why it benefits from using intra-layer representations for classification. We further demonstrate the advantages of leveraging cross-correlations between intra-layer and inter-layer features to decorrelate class-prototypes. (2) Building on our insights, we propose LayUP, a novel class-prototype approach to CL that leverages second-order statistics of features from multiple layers of a pre-trained model. Inspired by prior works [71, 40], we experiment with different methods for parameter-efficient model tuning and extend our method towards first session adaptation. Our final approach is conceptually simple, light in memory and compute, and works out-of-the-box with any transformer model. (3) We report performance improvements with a pre-trained Vit-B/16 [12] backbone on several benchmarks both in the Class-Incremental Learning (CIL) and in the more challenging Online Continual Learning (OCL) setting and show that LayUP is especially effective under large distributional shifts and in the low-data regime. Our results highlight the importance of taking a deeper look at the intermediate layers of pre-trained models to better leverage their powerful representations, thus identifying promising directions for CL in the era of large foundation models.
2 Related Work
2.1 Continual learning
Traditional approaches to CL consider training a model from scratch on a sequence of tasks, while striking a balance between adapting to the currently seen task and maintaining high performance on previous tasks. They can be broadly categorized into regularization [27, 67, 33, 1, 11], which selectively restrict parameter updates, (pseudo-)rehearsal [51, 4, 2, 47], which recover data either from a memory buffer or from a generative model, and dynamic architectures [53, 66, 55], which allocate disjoint parameter spaces for each task.
2.2 Continual learning with pre-trained models
Leveraging the powerful representations from large foundation models for downstream continual learning has shown to not only facilitate knowledge transfer, but also to increase robustness against forgetting [49, 43]. L2P [65] and DualPrompt [64] outperform traditional CL baselines by training a pool of learnable parameters (i.e., prompts) that are used as queries to guide a Vision Transformer (ViT) [12] towards adapting to downstream tasks. They serve as the basis for several more recent prompt learning methods such as S-Prompt [63], DAP [24], and CODA-Prompt [56], which further increase performance.
Contrary to prompt learning methods that keep the backbone parameters fixed while updating a small set of additional parameters, some parallel works explore methods for making careful adjustments to the backbone parameters. SLCA [69] uses a small learning rate for updates to the ViT backbone, while training a linear probing layer with a larger learning rate. L2 [57] applies regularization to the self-attention parameters of a ViT during continual fine-tuning. First Session Adaptation (FSA) [44] makes adjustments to the parameters of the backbone only during first task training.
As an alternative to the methods above that adjust the feature space via cross-entropy loss and backpropagation, constructing prototypes directly from the pre-trained features can avoid catastrophic forgetting during CL without storing data. Prior works [21, 72] find that a conceptually simple Nearest Mean Classifier (NMC) outperforms various prompt learning methods. ADAM [72] combines NMC with FSA and concatenates the features from the initial pre-trained ViT and the adapted parameters. Other methods leverage second-order feature statistics, i.e., covariance information, for class-prototype accumulation: FSA+LDA [44] applies an incremental version of linear discriminant analysis to a pre-trained ResNet encoder [14]. RanPAC [40] uses high-dimensional random feature projections to decorrelate class-prototypes of a pre-trained ViT backbone.
All the aforementioned approaches have in common that they consider only the last layer representations (or, features) for classification. We provide a different perspective to the prototype-based approach to CL and show that intra-layer representations can add valuable information to the linear transformation of class prototypes (cf. Eq. 5), which leads to better class separability and increased robustness to large distributional shifts.
3 Preliminaries
3.1 Continual learning problem setting
We consider a feature extractor $\phi(\cdot)$ composed of $L$ consecutive layers and a sequence of training datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$, where the $t$th task $\mathcal{D}_{t}=\{({\bm{x}}_{t,n},y_{t,n})\}_{n=1}^{N_{t}}$ contains pairs of input samples ${\bm{x}}_{t,n}\in\mathcal{X}_{t}$ and their ground-truth labels $y_{t,n}\in\mathcal{Y}_{t}$. Let $\mathcal{Y}_{t}$ denote the label space of the $t$th task and $\mathcal{Y}=\bigcup_{t=1}^{T}\mathcal{Y}_{t}$. The total number of classes seen in $\mathcal{D}$ (i.e., the number of unique elements in $\mathcal{Y}$) is denoted with $C$. We represent the $d_{l}$-dimensional encoded features from the $l$th layer of model $\phi$ from some test image ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$ with unknown label during inference as $\phi_{l}({\bm{x}})\in{\mathbb{R}}^{d_{l}}$. We consider rehearsal-free CL, where historical data can not be fetched for replay. Experiments in Sec. 5 are conducted under Class-Incremental Learning (CIL) and Online Continual Learning (OCL) settings which prohibit the access of the model to task identifiers during testing (notably, our method does not need any task-specific information during training, either). OCL is a more tightened definition of CIL towards each datum in the stream of tasks being observed only once.
3.2 Class-prototype methods for CL with pre-trained models
To leverage the powerful representations from large-scale pre-training while overcoming the limitations of full body adaptation and linear probing, class-prototype methods construct representatives for each class from the last layer features of a pre-trained model. The most straightforward class-prototype method is the Nearest Mean Classifier (NMC), which aggregates per-class training sample features via averaging (denoted as $\overline{{\bm{c}}}_{y}$ for each class $y$) and, during inference, selects for each test sample’s feature vector the class with the smallest Euclidean distance [21] or highest cosine similarity [72] with its class-prototype. Considering the cosine similarity measure and some ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, the predicted class label is obtained by:
$\hat{y}=\operatorname*{arg\,max}_{y\in\{1,\dots,C\}}s_{y},\quad s_{y}:=\frac{\phi_{L}({\bm{x}})^{T}\ \overline{{\bm{c}}}_{y}}{\lVert\phi_{L}({\bm{x}})\rVert\cdot\lVert\overline{{\bm{c}}}_{y}\rVert},$ (1)
[…]
$\hat{y}=\operatorname*{arg\,max}_{y\in\{1,\dots,C\}}s_{y},\quad s_{y}:=\phi_{L}({\bm{x}})^{T}\ ({\bm{G}}+\lambda{\bm{I}})^{-1}\ {\bm{c}}_{y}$ (2)
for a regression parameter $\lambda\geq 0$, $d_{L}$-dimensional identity matrix ${\bm{I}}$, and ${\bm{G}}$ expressed as summation over outer products as
${\bm{G}}=\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\phi_{L}({\bm{x}}_{t,n})\otimes\phi_{L}({\bm{x}}_{t,n})$ (3)
Although Eq. 1 and Eq. 2 are defined with respect to the maximum number of classes $C$ after observing all $T$ tasks, they can be applied after seeing any task $t\leq T$ or any sample $n\leq N_{t}$. When denoting the extracted features of some input datum ${\bm{x}}_{t,n}\in\mathcal{D}_{t}$ as ${\bm{f}}_{t,n}=\phi_{L}({\bm{x}}_{t,n})$ and considering ${\bm{F}}\in\mathbb{R}^{N\times d_{L}}$ as concatenated row-vector features ${\bm{f}}_{t,n}$ of all $N$ training samples, Eq. 3 will be reduced to ${\bm{G}}={\bm{F}}^{T}{\bm{F}}$, which corresponds to the original definition of a Gram matrix as used in the closed-form ridge estimator [18].
4 LayUP: Intra-Layer Universal Prototypes
4.1 Motivation
We argue that a combination of (i) enriching the last layer representations with hierarchical intra-layer features and (ii) decorrelating intra-layer features, which represent image properties such as content and style, via Gram matrix transformation increases robustness to domain shifts and thus improves generalizability to downstream continual tasks. To this end, we propose to integrate embeddings from multiple layers via concatenation into a ridge regression estimator as defined in Eq. 2. Let
$\Phi_{-k:}({\bm{x}})=\left(\phi_{L-k+1}({\bm{x}}),\dots,\phi_{L-1}({\bm{x}}),\phi_{L}({\bm{x}})\right)$ (4)
denote the concatenated input features of some input sample ${\bm{x}}\in\mathcal{D_{\mathrm{test}}}$, extracted from the last $k$ layers of the pre-trained model $\phi(\cdot)$. Such features can be, e.g., class embeddings ([CLS]) of a transformer [59] encoder or flattened feature maps of a ResNet [14] encoder. This yields a modified estimator
$s_{y}:=\Phi_{-k:}({\bm{x}})^{T}\ \left({\bm{G}}+\lambda{\bm{I}}\right)^{-1}\ {\bm{c}}_{y}$ (5)
with $(d_{L-k+1}+\dots+d_{L-1}+d_{L})$-dimensional identity matrix ${\bm{I}}$ and
${\bm{G}}=\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\Phi_{-k:}({\bm{x}}_{t,n})\otimes\Phi_{-k:}({\bm{x}}_{t,n})$ (6)
We first motivate our approach by showing that intra-layer representations capture expressive class statistics for prototype-based classification. For the split CIFAR-100 ($C=100$) and ImageNet-R ($C=200$) datasets, we construct one prototype-based classifier using Eq. 2 per layer $l\in\{1,\dots,L\}$ and count the total number of classes that layer $l$ predicts better than any other layer $l^{\prime}\neq l$. Results are provided in Fig. 1(a). We observe that especially the last four intermediate layers (8–11) account for a top performance of between 25% and 54% of all classes, indicating that the representations obtained by intermediate layers can generally add meaningful knowledge to better separate classes at different hierarchical levels.
(a) Intra-layer performance
(b) LayUP (shared) vs. averaging over layer-wise similarity scores (separate)
Figure 2: (a) Comparison of classification performance of layers $(1,\dots,L)$ for two different pre-training paradigms of a ViT-B/16 [12] backbone and split ImageNet-R and CIFAR-100 datasets. Each bar indicates the absolute count of classes for which a classifier constructed using Eq. 2 for the respective layer yields highest classification accuracy compared to all other layers.
(b) Average accuracy (%) over CL training on split ImageNet-R and CIFAR-100 datasets following phase B of Alg. 1: Comparison of LayUP implementations for different values of $k$ using Eq. 5 versus averaging over separate similarity scores for each layer using Eq. 2.
There are two intuitive ways to integrate intra-layer representations into the class-prototype classifier: Averaging over $k$ separate classifiers per Eq. 2 or concatenating representations from the last $k$ layers to obtain shared class prototypes per Eq. 5. Fig. 1(b) shows average accuracies for two split datasets and different values of $k$. For each $k$, using shared class-specific information as per Eq. 5 is superior to averaging over separate classifiers, showing that the correlations across layers, which are captured in the shared Gram matrix (cf. Eq. 6), add meaningful information to enhance class separability and thus improve performance. Although this leads us to choose higher memory demands compared with a last layer classifier and averaging over layer-wise classifiers, we will show in Sec. 4.3 that our method is still considerably lighter in memory and computation than other competitive class-prototype methods for CL.
Interestingly, the performance of the last layer classifier is roughly the same compared with averaging over separate classifiers for all choices of $k$. This observation refutes the argument that class embeddings from multiple consecutive layers that are individually strong at predicting specific classes might, when combined, introduce noise that impairs performance, which is clearly not the case in Fig. 1(b).
4.2 Combination with parameter-efficient model adaptation
A benefit of class-prototype methods for CL is that they can be orthogonally combined with adaptation techniques to refine the pre-trained representations. As accumulated class prototypes are subject to discrepancy upon distributional shifts during supervised fine-tuning, any adaptation strategies to the latent representations should be carefully tailored to the class-prototype method used. Previous works [44, 72] have found adaptation to downstream continual tasks during First Session Adaptation (FSA) on $\mathcal{D}_{1}$ as sufficient to bridge the domain gap while maintaining full compatibility with CL.
To maintain the powerful generic features of the pre-trained backbone, we choose to apply FSA to additional parameter-efficient transfer learning (PETL) parameters while keeping the backbone frozen throughout. To update PETL parameters during FSA, we train a linear classification head with $N_{1}$ outputs via Adam optimizer and cross-entropy loss and discard the classification head afterward. During the CL phase, we keep all model parameters frozen and update ${\bm{G}}$ and ${\bm{c}}_{y}$ only.
We follow prior works [72, 40] and experiment with Visual Prompt Tuning (VPT) [22], Scaling and Shifting of Features (SSF) [34], and AdaptFormer [5] as PETL methods and refer to the cited papers for details. The pseudocode of the LayUP algorithm for the CIL setting is depicted in Alg. 1. For the OCL setting, the FSA stage is omitted and a fixed $\lambda$ is used, as a dynamic $\lambda$ search requires multiple passes over the input data.
Algorithm 1 LayUP Training
pre-trained network $\phi(\cdot)$PETL parameters data $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{T}\}$$k$
[…]
Collect $\phi_{L}({\bm{x}})$
[…]
Collect $\Phi_{-k:}({\bm{x}}_{t})$ (Eq. 4)
${\bm{G}}\leftarrow{\bm{G}}+\Phi_{-k:}({\bm{x}}_{t})\otimes\Phi_{-k:}({\bm{x}}_{t})$
${\bm{c}}_{y_{t}}\leftarrow{\bm{c}}_{y_{t}}+\Phi_{-k:}({\bm{x}}_{t})$EndFor
# Cf. Sec. A.1
Optimize $\lambda$ to compute $({\bm{G}}+\lambda{\bm{I}})^{-1}$EndFor
[…]
4.3 Memory and runtime complexity comparisons
We compare the memory and runtime requirements of LayUP with the three competitive CL methods ADAM [72], SLCA [69], and RanPAC [40] for a typical class count of $C=200$. We neglect the memory cost of the ViT backbone, the fully connected classification layer, and the class prototypes, as they are similar to all baselines. LayUP stores $\sim$1M PETL parameters (AdaptFormer [5]) and an additional intra-layer Gram matrix in memory. The size of this layer depends on the choice of $k$ and has $(d_{L-k+1}+\dots+d_{L-1}+d_{L})^{2}$ entries (note that for ViT-B/16 [12] model, every layer outputs tokens of the same dimensionality, such that $d_{l}=768$ for $l\in\{1,\dots,L\}$). For $k=6$, which we use in our final experiments, the matrix stores $\sim$21M values. RanPAC updates a Gram matrix of size 100M and requires an additional $\sim$11M parameters for the random projection layer and PETL parameters. SLCA stores $C$ additional covariance matrices of size $d_{L}^{2}$ with a total of $\sim$118M entries. Finally, ADAM stores a second copy of the feature extractor, necessitating 84M additional parameters. Assuming a comparable memory cost for model parameters and matrix entries, our method reduces additional memory requirements by $81\%$, $82\%$, and $75\%$ compared with RanPAC, SLCA, and ADAM, respectively.
[…]
5 Experiments
In what follows, we conduct a series of experiments to investigate the performance of our approach under different CL settings. First, we provide an overview of the datasets and implementation details in Sec. 5.1. Second, we perform empirical comparisons of our method with strong recent baselines under the CIL and the more challenging OCL settings in Sec. 5.2 and Sec. 5.3, respectively. Third, we conduct ablation studies in Sec. 5.4. We conclude our empirical analysis by investigating the contribution of different choices of $k$ last layers to the classifier performance in Sec. 5.5.
5.1 Datasets and implementation details
Method CIFAR IN-R IN-A CUB OB VTAB Cars
Joint full FT $93.6$ $86.6$ $71.0$ $91.1$ $80.3$ $95.5$ $83.7$
Joint linear probe $87.9$ $71.2$ $56.4$ $89.1$ $78.8$ $90.4$ $66.4$
L2P [65] $84.6$ $72.5$ $42.5$ $65.2$ $64.7$ $77.1$ $38.2$
DualPrompt [64] $81.3$ $71.0$ $45.4$ $68.5$ $65.5$ $81.2$ $40.1$
CODA-P [56] $86.3$ $75.5$
ADAM [72] $87.6$ $72.3$ $52.6$ $87.1$ $74.3$ $84.3$ $41.4$
SLCA [69] $91.5$ $77.0$ $59.8^{*}$ $84.7$ $73.1^{*}$ $89.2^{*}$ $67.7$
RanPAC [40] $78.1$ $61.8$ $92.6$ $77.7$
LayUP $91.5$ $88.1$ $77.6$
Ablations
w/o FSA $88.7$ $73.4$ $51.6$ $86.9$ $77.0$ $92.9$ $77.5$
${k=1}$ $90.4$ $79.0$ $61.5$ $86.0$ $74.3$ $93.1$ $75.5$
$k=1$, w/o FSA $86.5$ $69.2$ $55.6$ $85.4$ $72.8$ $92.2$ $69.2$
LayNMC $88.1$ $75.2$ $56.2$ $85.5$ $71.8$ $87.5$ $57.6$
Table 1: Average accuracy (%) after training: Comparison of prompting, backbone fine-tuning, and class-prototype methods for the CIL setting. Results are taken from [40] except results for SLCA and Coda-P, which are taken from the respective papers. Results for SLCA marked with ($*{}^{*}$) were reproduced using the officially released code. The remaining results for CODA-P could not be reproduced due to excessive GPU memory requirements for training ($>$24GB).
Following prior works [64, 65, 72, 69, 40], we experiment with two ViT-B/16 [12] models, the former with self-supervised pre-training on ImageNet-21K (ViT-B/16-IN21K) [52], the latter with additional supervised fine-tuning on ImageNet-1K (ViT-B/16-IN1K) [30]. During adaptation in the CIL setting, PETL parameters are trained using a batch size of 48 for 20 epochs and Adam [26] with momentum for optimization. We employ learning rate scheduling via cosine annealing, starting with 0.03. We train five prompt tokens for VPT and use a bottleneck dimension of 16 for AdaptFormer. All baselines use the same aforementioned ViT-B/16 backbones and are trained on seven representative split datasets CIFAR-100 (CIFAR) [29], ImageNet-R (IN-R) [16], ImageNet-A (IN-A) [17], CUB-200 (CUB) [60], OmniBenchmark (OB) [70], Visual Task Adaptation Benchmark (VTAB) [68], and Stanford Cars-196 (Cars) [28]. None of the methods compared in Tab. 1 and Tab. 2 require a data buffer for rehearsal. We use $T=10$ for all datasets, except $T=5$ for VTAB, which is a benchmark composed of five different datasets to be learned consecutively. (We report additional results for $T=5$ and $T=20$ in Sec. C.4.) VTAB can be considered as a hybrid class- and domain-incremental learning benchmark according to the definition in [58], as label spaces of each task are partially overlapping. Dataset statistics can be found in Appendix B. For comparison, we use the average accuracy [36] metric $A_{t}=\frac{1}{t}\sum_{i=1}^{t}R_{t,i}$, with $R_{t,i}$ denoting the classification accuracy on the $i$th task after training on the $t$th task. All experiments are conducted with $k=6$ unless stated otherwise. Final accuracy $A_{T}$ after learning the last task is reported for random seed 1993 and the best combination of PETL method and ViT backbone (similar to [40, 72]) in the main paper. We refer to Sec. C.1 for analysis of each $A_{t}$, forgetting measures, and variability across random initialization, and to Sec. C.3 for average accuracy over training for different PETL methods and pre-trained backbones in the CIL setting.
5.2 Performance in the CIL setting
Method CIFAR IN-R IN-A CUB OB VTAB Cars
Sequential FT $12.8$ $5.3$ $6.9$ $5.1$ $3.3$ $7.3$ $10.9$
NMC [72] $83.4$ $61.2$ $49.3$ $86.7$ $73.2$ $88.4$ $37.9$
[…]
Ablations
${k=1},\lambda=0$ $86.5$ $69.6$ $85.4$ $72.8$ $92.2$ $69.2$
${k=1},\lambda=1$ $86.5$ $69.6$ $85.4$ $76.3$ $92.3$ $69.2$
LayNMC $80.8$ $59.7$ $50.4$ $82.9$ $69.3$ $85.9$ $40.0$
Table 2: Average accuracy (%) after training: Comparison of class-prototype methods for the OCL setting. Results for RanPAC and NMC are reproduced according to the officially released code in [40].
We first compare LayUP with several prompt learning, fine-tuning, and class-prototype methods for CL. We additionally report results for joint training of a randomly initialized linear probe and joint full fine-tuning (FT) of the ViT-B/16 backbone. Final accuracy scores are shown in Tab. 1.
LayUP surpasses all baselines for four of the seven split datasets. Interestingly, these four datasets (IN-R, IN-A, VTAB, Cars) have two distinctive characteristics: First, they have the highest domain gap (indicated by the cosine distance to the pre-trained miniImageNet domain for the offline setting and 50 shots, cf. [44]) among all datasets. Second, they have significantly less training data than CIFAR, OB, and CUB (cf. Appendix B). The former confirms our initial hypothesis that intra-layer representations are more domain-invariant and consequently more robust to large distributional shifts. The latter indicates that especially in the low-data regime, in contrast to other methods that tend to overfit to the target domain, LayUP constructs class prototypes and decision boundaries that generalize well even if the amount of training data is scarce compared with the data used for ViT pre-training. It is further noteworthy that RanPAC, which is the only baseline that LayUP does not consistently outperform, is considerably more expensive regarding both memory and computation (cf. Sec. 4.3).
5.3 Performance in the OCL setting
We are interested in assessing the performance of our method in the challenging OCL setting, where only a single pass over the continual data stream is allowed. All baselines are class-prototype methods for continual learning on a frozen embedding network (thus we omit FSA stages for RanPAC, and our approach), except for sequential fine-tuning, where we update the parameters of the pre-trained ViT via cross-entropy loss and Adam [26] optimizer during a single epoch. As the choice of ridge regression [18] parameter as used in RanPAC and our method requires prior knowledge about the downstream continual data–which we do not have in a realistic streaming learning setting–we compare with two simplified versions of the Gram matrix inversion $({\bm{G}}+\lambda{\bm{I}})^{-1}$ (cf. Eq. 5 and Eq. 2): In the first variant, we omit $\lambda$-based regularization completely, such that $\lambda=0$ and the Gram matrix is inverted without regularization (i.e., ${\bm{G}}^{-1}$). In the second variant, e.g., as used in [44], we compare with $\lambda=1$, such that we linearly transform identity-regularized class prototypes (i.e., $({\bm{G}}+{\bm{I}})^{-1}$).
As indicated in Tab. 2, unrestricted fine-tuning of the backbone is detrimental to the generalizability of the pre-trained embeddings and leads to forgetting and low performance. Consequently, the sequential FT baseline is outperformed by class-prototype methods in the OCL setting for most benchmarks. LayUP is largely robust to missing regularization ($\lambda=0$), with the exception of a drop in performance on VTAB. Moreover, it maintains a high performance across all benchmarks for regularization with $\lambda=1$. On the contrary, RanPAC shows high variability in performance across benchmarks under the absence of regularization and for regularization with $\lambda=1$, partly resulting in near-zero accuracy. Such results can be explained by the fact that the high-dimensional Gram matrix obtained after random projections in RanPAC overfits the training data and thus struggles to generalize, especially under a complete lack of regularization. The higher overall robustness to different OCL settings makes LayUP a favorable approach in continual learning scenarios where the length of the input stream and the nature of the data might not be known in advance.
5.4 Ablation Study
[…]
LayUP consistently benefits from intra-layer representations for class-prototype construction, with the sole exception of training on IN-A in the OCL setting (Tab. 2). It further benefits from leveraging second-order statistics for the decorrelation of class prototypes and first session training to bridge the domain gap to downstream CL tasks. Finally, it is the combination of (i), (ii), and (iii) that yields the highest performance and thus makes LayUP a strong and competitive class-prototype method for CL with pre-trained models.
5.5 Analysis of intra-layer representation depth
To provide more insights into the impact of the intra-layer representation depth on LayUP performance, we plot in Tab. 3 the average accuracy after training for different choices of $k$ using a Vit-B/16-IN1K backbone and AdaptFormer [5] as PETL method.
$k$ CIFAR IN-R IN-A CUB OB VTAB Cars
[…]
Table 3: Average accuracy (%) after training: LayUP performance for different values of $k$ for a pre-trained Vit-B/16-IN1K and FSA with AdaptFormer. The 1st, 2nd, and 3rd highest scores and the 1st, 2nd, and 3rd lowest scores are highlighted.
We observe that LayUP configurations with $k\geq 4$ consistently surpass the final layer classifier (i.e., $k=1$), which shows that class-prototype methods for CL can generally benefit from intra-layer representations. Although the optimal choice of $k$ differs for each dataset, the results are indicative of a choice of larger $k$. As a larger value of $k$ directly corresponds to increased memory and runtime complexity during inference, its choice should be carefully weighed up, taking into account demands and restrictions at different dimensions.
5.6 How universal are intra-layer prototypes?
We are interested in determining how universal the intra-layer prototypes are for classification, i.e., how broad the range of classes is that benefit from LayUP. For this purpose, we run phase B of Alg. 1 and provide the fraction of classes for each dataset that has an increased or equal final accuracy score with LayUP for $k=6$ compared with LayUP for $k=1$ (last layer Gram-inverted classifier, cf. Eq. 2).
CIFAR IN-R IN-A CUB OB VTAB Cars
$94$% $79$% $81$% $89$% $80$% $94$% $83$%
Table 4: Percentage of classes per dataset that are better or equally classified with LayUP for $k=6$ compared with the last layer Gram-inverted classifier (Eq. 2).
Tab. 4 shows that between 80% and 94% of the total classes in each dataset are better or equally classified upon introducing intermediate representations to class-prototype construction, indicating the universal applicability of intra-layer prototypes to a broad range of tasks, classes, or domains.
6 Conclusion
In this paper, we propose LayUP, a simple yet effective rehearsal-free class-prototype method for continual learning with pre-trained models. LayUP leverages intra-layer representations of a pre-trained feature extractor to increase robustness and generalizability, especially under large domain gaps and in low-data regimes. It further computes second-order feature statistics to decorrelate class prototypes, combined with parameter-efficient first-task adaptation. Extensive experiments across a range of image classification datasets and continual learning settings demonstrate that LayUP performs strongly against competitive baselines at a small fraction of their memory and compute cost. In future work, we will further investigate integrating class-prototype methods with continual adaptation of pre-trained models beyond first-task adaptation, which is particularly beneficial for learning under multiple distributional shifts.
Supplementary Material
Appendix A Training and Implementation Details
A.1 Optimization of ridge regression parameter $\lambda$
We perform the optimization over the regression parameter $\lambda$ (cf. Alg. 1) as follows: For every task $t$, we perform a random 80:20 split of the training data stratified by ground truth labels. We then update ${\bm{G}}$ and ${\bm{c}}_{y}\ \forall\ y\in\mathcal{Y}_{t}$ (which were already updated during task $11,\dots,t-$ 1 training) for the first 80% of training data of task $t$ and then choose the value of $\lambda\in\{10^{-3},10^{-2},10^{-1},10^{0},10^{1},10^{2},10^{3}\}$ that yields highest per-sample accuracy on the remaining 20% of training data. Finally, we update ${\bm{G}}$ and ${\bm{c}}_{y}\ \forall\ y\in\mathcal{Y}_{t}$ for those 20% of training data of task $t$ and repeat the same process with $t+1$.
A.2 Data augmentation
During the training process, data augmentation was applied to all datasets, incorporating random cropping of the images to varying sizes, ranging from 70% to 100% of their original dimensions, while maintaining an aspect ratio between 3:4 and 4:3. After the resizing, images were randomly flipped horizontally and brightness, contrast, saturation, and hue were varied randomly in a 10% range. Finally, the images were center-cropped to 224×224 pixels for all datasets, except for CIFAR-100, where images were directly resized from the original 32×32 to 224×224 pixels. During inference, images of all datasets were resized to 224x224 pixels without further modification.
A.3 Compute resources
All experiments in this work were conducted on an Ubuntu system version 20.04.6 with a single NVIDIA GeForce RTX 3080 Ti (12GB memory) GPU.
Appendix B Datasets
We provide a summary of the datasets compared in the main experiments in Tab. 5. CIFAR-100 (CIFAR) contains 100 classes of natural images of different domains and topics and can be considered relatively in-distribution with the pre-train domain of ImageNet-1K and ImageNet-21K. ImageNet-R (IN-R) contains image categories overlapping with ImageNet-1K, but is a selection of out-of-distribution samples for the pre-train dataset that are either hard examples or newly collected data of different styles. ImageNet-A (IN-A) likewise has overlapped categories with ImageNet-1K, but comprises real-world adversarially filtered images that fool existing ImageNet pre-trained classifiers. The Caltech-UCSD Birds-200-2011 (CUB) dataset is a specialized collection of labeled images of 200 bird species, encompassing a diverse range of poses and backgrounds. OmniBenchmark (OB) serves as a compact benchmark designed to assess the generalization capabilities of pre-trained models across semantic super-concepts or realms, encompassing images of 300 categories that represent distinct concepts. The Visual Task Adaptation Benchmark (VTAB) as used in our work is a composition of the five datasets Resisc45 [8], DTD [10], Pets [46], EuroSAT [15], and Flowers [42], that are learned consecutively to emulate the emergence of different domains. The Stanford Cars-196 (Cars) dataset comprises images of cars belonging to one of 196 unique combinations of model and make.
Original CL $T$ $N_{\text{train}}$ $N_{\text{val}}$ $C$
CIFAR [29] [51] 10 50 000 10 000 100
IN-R [16] [64] 10 24 000 6 000 200
IN-A [17] [72] 10 6 056 1 419 200
CUB [60] [72] 10 9 465 2 323 200
OB [70] [72] 10 89 668 5 983 300
VTAB [68] [72] 5 1 796 8 619 50
Cars [28] [69] 10 8 144 8 041 196
Table 5: Overview of datasets: Original publication, formulation for the CL setting, number of tasks ($T$) in the main experiments in Sec. 5, number of training samples ($N_{\text{train}}$), number of validation samples ($N_{\text{val}}$), and number of classes ($C$).
Appendix C Additional Results
We use the following two metrics in our work for experimental evaluation: Average accuracy $A_{t}$ (following the definition of [36]) and average forgetting $F_{t}$ (following the definition of [3]). Average accuracy is defined as
$A_{t}=\frac{1}{t}\sum_{i=1}^{t}R_{t,i},$ (7)
where $R_{t,i}$ denotes the classification accuracy on task $i$ after training on task $t$. Using the same notion of $R_{t,i}$, average forgetting is defined as
$F_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}\operatorname*{arg\,max}_{t^{\prime}\in\{1,\dots,t-1\}}R_{t^{\prime},i}-R_{t,i}$ (8)
As the number of different classes to choose from in a CIL and an OCL setting grows as new tasks are being introduced, $A_{t}$ tends to decrease and $F_{t}$ tends to rise over the course of training. While we only report on average accuracy metric in the main paper, the experimental results in Sec. C.1, Sec. C.2, and Sec. C.3 are reported both with respect to average accuracy and average forgetting.
C.1 Variability across seeds and performance over time
LayUP accuracy and forgetting for all seven datasets over the course of class-incremental training (Phase B in Alg. 1) is depicted in Fig. 3. There are almost no differences between final results after the last task across seeds, which indicates LayUP being insensitive to different task orders and initialization of PETL parameters. The highest variability can be observed for the VTAB benchmark, where each task represents a completely new domain. Therefore, the task order–and consequently the domain that the representations are adjusted to during first session adaptation–has a greater influence on the generalization capabilities of the model during CL. As expected, the average forgetting increases as more tasks are being introduced, as the model has more classes to tell apart during inference.
Figure 3: Average accuracy (left) and average forgetting (right) after training on each task $t$ in the CIL setting: Variability across random seeds for each ViT-B/16 models after first session training with AdaptFormer as PETL method. Results are reported for seeds 1993-1997 to ensure reproducibility with the resulting standard error indicated by shaded area.
C.2 Experiments with different layer choice $k$
To analyze the learning behavior over time in the CIL setting among different choices of maximum representation depth $k$ for prototype construction, we plot average accuracy and forgetting for $k=1$ (last layer only), $k=6$, and $k=12$ (all layers) in Fig. 4. Clearly, the classification performance based on last layer representations only is inferior to intra-layer representations as used in LayUP, as it leads to both a lower accuracy and a higher forgetting rate. At the same time, there is no difference in performance between $k=6$ and $k=12$ evident, indicating that early layers of the model do not add meaningful knowledge to the classification. Given that the choice of $k$ is subject to a trade-off between performance gain and memory and computational cost, the results confirm $k=6$ as used in the main experiments in Sec. 5 to be a reasonable choice.
Figure 4: Average accuracy (left) and average forgetting (right) after training on each task $t$ in the CIL setting: Comparison of different choices of $k$ last layers for prototype construction ($k\in\{1,6,12\}$) and ViT-B/16 models after first session training with AdaptFormer as PETL method.
C.3 Experiments with different backbones and PETL methods
Following prior works [72, 40], we experiment with different PETL methods and ViT-B/16 models for LayUP, as the additional benefit from additional fine-tuning of the representations differs depending on the characteristics of the downstream domain [44]. Results are shown in Fig. 5.
In all but one dataset, we found adapter methods (AdaptFormer) to be superior to be superior to prompt learning (VPT) or feature modulation (SSF). However, there is no combination of PETL method and pre-trained model that generally outperforms all other variants. Such results confirm the findings of [44] and underline the importance of considering different pre-training schemes and strategies for additional fine-tuning.
Figure 5: Average accuracy (left) and average forgetting (right) after training on each task $t$ in the CIL setting: Comparison of different PETL methods (AdaptFormer [5], SSF [34], and VPT [22]) and ViT-B/16 models.
C.4 Experiments with different task counts $T$
To show whether a benefit of intra-layer representations can be observed for different task counts, we compare average accuracy scores after training for six different datasets, three different task counts ($T\in\{5,10,20\}$) and three different choices of $k$ last layers for class-prototype generation. $k=1$ corresponds to classification only based on the last layer (i.e., final) representations of the backbone, as it is done in prior work. $k=6$ means to concatenate layer-wise features from the latter half of the network layers. Finally, $k=12$ uses concatenated features of all layers of the pre-trained ViT for classification. Results are presented in Tab. 6.
Performance scores across task counts and datasets are consistently higher for intra-layer representations ($k=6$ and $k=12$) compared with last-layer-only representations ($k=1$). However, differences in performance between $k=6$ and $k=12$ are not apparent (with the exception of a $T=20$ split on the CUB dataset, which can be attributed to some sensitivity to initialization of adapter parameters). Considering that a higher $k$ corresponds to an increased computational and memory demand as ${\bm{G}}$ and ${\bm{c}}_{y}$ increase in dimension, this confirms $k=6$ as reasonable choice for maximum representation depth.
$k$ $T$ CIFAR IN-R IN-A CUB OB Cars
1 5 89.9 80.6 62.5 85.3 73.2 75.6
10 89.1 79.1 61.4 85.5 71.6 75.6
20 86.3 77.8 59.5 83.5 72.2 75.5
6 5 91.3 83.0 64.5 87.6 78.0 82.2
10 90.8 81.9 62.5 87.7 76.6 81.8
20 90.3 80.4 61.6 69.5 77.8 81.8
12 5 91.6 82.7 62.7 87.1 77.6 81.8
10 91.1 82.1 62.7 87.5 73.8 81.7
20 90.5 80.3 56.8 86.0 78.3 81.9
Table 6: Average accuracy (%) after training: Comparison of different task counts $T$ for $k=1$ (prototype construction from last layer only), $k=6$ (as used in Sec. 5), and $k=12$ (prototype construction from all network layers). Scores listed are for AdaptFormer and ViT-B/16-IN1K. VTAB has a fixed number of datasets that are treated as tasks (thus $T=5$) and is therefore omitted in the comparison.